Hello Mr. McNeal we are happy to provide consulting services to you here at Nelson Consulting Service. It is my responsibility to ensure you receive accurate and clear consulting with your requests. I understand you are trying to maximize your profits in your fund raiser. To be clear you have a budget of $4000 and you want to buy and sell wigs, ball and a portable charger. You stated, wigs will cost $10 to buy and will be sold for $30, Balls will cost $3 to buy and will be sold for $12 and the Chargers will cost $10 to buy and sold for $24. Additionally you have a table which is 60x30 in coming to a total of 1800 sqin. Your items take up 10 sqin for wigs, 6 sqin for balls and 2 sqin for chargers. To Maximize the profits out the 3 items which 2 should be bought and sold?  How many of the items can be sold using the Simplex Method? Show work.
Answer #1

Estimated budget = $4000

Estimated profit:

Wig :

Cost = $10

Sales = $30 Hence Profit = 30-10 = 10 per unit

Balls:

Cost $3

Sales $12 Hence profit = 12-3 = 9 per unit

Charger:

Cost $10

Sales $24 Hence profit = 24-10 = $14 per unit

Producing charger and selling them incurs higher profits.

Hence wigs and balls can be bought and sold while charger can be produced as it incurs less production cost and gives higher profit